Change page style: 

NIRI Update

24 September 2002

Over the weekend NIRI's pupil wheel failed, making further science and engineering work with NIRI impossible. We are warming NIRI now, and will fix the problem as quickly as we can. At present, queue observations have been suspended. We will know within the next week or two what will be required to get NIRI working reliably and back on the telescope. We will provide an update on the NIRI web pages at that time. NIRI PIs should continue to prepare for observations, including submitting and modifying Phase II programs as necessary.

Two weeks ago we found a problem with the drive mechanism for the on-instrument wavefront sensor gimbal mirror. This problem did not impact our ability to execute the science programs in the queue because we can also use the peripheral wavefront sensors. Now that NIRI will be warm and open to fix the pupil wheel, we will also fix the gimbal mirror in preparation for science use and Altair commissioning later this year.

It was reassuring to see NIRI performing well prior to the pupil wheel failure. For three nights of good weather, the telescope and NIRI consistently delivered 0.3 arcsec images at 2.2 microns. In that short time, data were taken for four programs, two of which were completed. We look forward to completing more programs this semester.


Last update Sept. 24, 2002; Joe Jensen